Find the Im(w)where w is the positive solution to z2 =6+5i
Answer #1

2 find The Im (w), where wich Politive solution to 7 = 6455 - Se 74² = 1675 - 136 + 25 = 561 5 -0.695 6 O-ran (y) = rant z r